# Methodology for an Analysis of Influencing Factors on 3D Object Detection Performance thanks: The research is accomplished within the project “AUTOtech.agil” (FKZ 01IS22088S). We acknowledge the financial support for the project by the Federal Ministry of Education and Research of Germany (BMBF)

### v2

@@ Abstract
In automated driving, object detection is an essential task to perceive the environment by localizing and classifying objects. Most object detection algorithms are based on deep learning for superior performance. However, their black-box nature makes it challenging to ensure safety. In this paper, we propose a first-of-its-kind methodology for analyzing the influence of various factors related to the objects or the environment on the detection performance of both LiDAR- and camera-based 3D object detectors. We conduct a statistical univariate analysis between each factor and the detection error on pedestrians to compare their strength of influence. In addition to univariate analysis, we employ a Random Forest (RF) model to predict the errors of specific detectors based on the provided meta-information. To interpret the predictions of the RF and assess the importance of individual features, we compute Shapley Values. By considering feature dependencies, the RF captures more complex relationships between meta-information and detection errors, allowing a more nuanced analysis of the factors contributing to the observed errors. Recognizing the factors that influence detection performance helps identify performance insufficiencies in the trained object detector and supports the safe development of object detection systems.
I Introduction
@@ II Related Work
III Methodology
This section describes the overall methodology used to investigate the factors that influence detection performance. First, we present the meta-information and the selected object detectors. We then outline the methodology our analysis.
III-A Meta-information
@@ IV Results
IV-A False Negative Errors
@@
Differences between the LiDAR-based detector are also observed. As observed in the Kendall’s Tau analysis in Fig. 1, the MI-analysis in Fig. 3 shows that PointPillars is more sensitive to pedestrian attributes and velocity. In contrast, PointPillars appears to be less sensitive to the visibility token respectively occlusion compared to Megvii and Transfusion-L, a behavior not observed before in the Kendall’s Tau analysis in Fig. 1.
Following the univariate analysis, Fig. 3 shows the results of the meta-model-based analysis. As previously described, an RF model is trained to predict whether an existing pedestrian is detected (TP) or not (FN) based on meta-information. In order to quantify the importance of each meta-information variable, the mean absolute SHAP values are computed to assess their impact on the RF model’s prediction. Consistent with the results from Kendall’s Tau and MI analysis, the mean absolute SHAP values in Fig. 3 indicate that distance, angular sizes, and the visibility token or the occlusion of the object have the strongest influence on the performance of the camera-based detectors. However, the vertical angular size has a greater effect than the horizontal angular size. While there are minor differences between the camera-based detectors, their general trends remain similar. In contrast, the influence of meta-information on LiDAR detectors is very different from the results of the univariate analysis in Fig. 1 and Fig. 3. Megvii and Transfusion-L show a similar behavior, expressed by a strong influence of occlusion, velocity, and the pedestrian’s attribute. These patterns are also observed in the MI-analysis in Fig. 3. However, unlike in the MI analysis, angular sizes play an important role in both detectors’ performance, as observed in Fig. 3. PointPillars, on the other hand, behaves quite differently from Megvii and Transfusion-L. While the ground-truth bounding box parameters generally have a low mean absolute SHAP value for all detectors, pedestrian length emerges as the most influential factor for PointPillars. Moreover, the key influencing factors of PointPillars, expressed by the mean absolute SHAP values in Fig. 3, also differ strongly from the previous MI-analysis in Fig. 3. The attribute, the velocity, and the occlusion, which have high MI-values, have relatively low SHAP values for PointPillars. Instead, in addition to the length of the pedestrian, the distance and the location have a large impact on the detection performance. In addition, relative wind direction and dew point temperature, factors not identified as influential in previous analyses, have a moderate influence on the detection performance of PointPillars.
@@
IV-B False Positive Error

### v3

@@ Abstract
In automated driving, object detection is crucial for perceiving the environment. Although deep learning-based detectors offer high performance, their black-box nature complicates safety assurance. We propose a novel methodology to analyze how object- and environment-related factors affect LiDAR- and camera-based 3D object detectors. A statistical univariate analysis relates each factor to pedestrian detection errors. Additionally, a Random Forest (RF) model predicts errors from meta-information, with Shapley Values interpreting feature importance. By capturing feature dependencies, the RF enables a nuanced analysis of detection errors. Understanding these factors reveals detector performance gaps and supports safer object detection system development.
1 Introduction
@@ 2 Related Work
3 Methodology
This section outlines the methodology for analyzing factors affecting detection performance, including meta-information, selected object detectors, and our analytical approach.
3.1 Meta-information
@@ 4 Results
4.1 False Negative Errors
@@
Differences between the LiDAR-based detector are also observed. As observed in the Kendall’s Tau analysis in Fig. 1, the MI-analysis in Fig. 3 shows that PointPillars is more sensitive to pedestrian attributes and velocity. In contrast, PointPillars appears to be less sensitive to the visibility token respectively occlusion compared to Megvii and Transfusion-L, a behavior not observed before in the Kendall’s Tau analysis in Fig. 1. Following the univariate analysis, Fig. 3 shows the results of the meta-model-based analysis. As previously described, an RF model is trained to predict whether an existing pedestrian is detected (TP) or not (FN) based on meta-information. In order to quantify the importance of each meta-information variable, the mean absolute SHAP values are computed to assess their impact on the RF model’s prediction. Consistent with the results from Kendall’s Tau and MI analysis, the mean absolute SHAP values in Fig. 3 indicate that distance, angular sizes, and the visibility token or the occlusion of the object have the strongest influence on the performance of the camera-based detectors. However, the vertical angular size has a greater effect than the horizontal angular size. While there are minor differences between the camera-based detectors, their general trends remain similar. In contrast, the influence of meta-information on LiDAR detectors is very different from the results of the univariate analysis in Fig. 1 and Fig. 3. Megvii and Transfusion-L show a similar behavior, expressed by a strong influence of occlusion, velocity, and the pedestrian’s attribute. These patterns are also observed in the MI-analysis in Fig. 3. However, unlike in the MI analysis, angular sizes play an important role in both detectors’ performance, as observed in Fig. 3. PointPillars, on the other hand, behaves quite differently from Megvii and Transfusion-L. While the ground-truth bounding box parameters generally have a low mean absolute SHAP value for all detectors, pedestrian length emerges as the most influential factor for PointPillars. Moreover, the key influencing factors of PointPillars, expressed by the mean absolute SHAP values in Fig. 3, also differ strongly from the previous MI-analysis in Fig. 3. The attribute, the velocity, and the occlusion, which have high MI-values, have relatively low SHAP values for PointPillars. Instead, in addition to the length of the pedestrian, the distance and the location have a large impact on the detection performance. In addition, relative wind direction and dew point temperature, factors not identified as influential in previous analyses, have a moderate influence on the detection performance of PointPillars.
@@
4.2 False Positive Error